OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Solve Log(2)3 + Log(2)x=3

OpenStudy (magbak):

@Hero @primeralph @satellite73

OpenStudy (magbak):

@thomaster

OpenStudy (magbak):

@countonme123 @Courage.

OpenStudy (anonymous):

Is the 2 in both logs a subscript?

OpenStudy (magbak):

YEs

OpenStudy (anonymous):

\[\log _{2}3+\log _{2}x=3\] That's how the equation looks right?

OpenStudy (magbak):

Yes

OpenStudy (magbak):

My time limit is about to end please help me.

OpenStudy (anonymous):

You should end up with \[\log _{2}3x=3\] Do you have answer options?

OpenStudy (magbak):

NO.

OpenStudy (magbak):

Just solve it. PLEASE I HAVE ONLY @ MINUTERS TO FINISH QUESTION PLEASE HELP <E/

OpenStudy (magbak):

<_- ->

OpenStudy (anonymous):

\[\log _{2}x=1\]

OpenStudy (magbak):

Thanks you.

OpenStudy (anonymous):

So x=2

OpenStudy (magbak):

Can you solve this one also\[8=3^(2x+2)\]

OpenStudy (magbak):

PLEASE I KNOW I AM PUTING PRESSURE BUT PLEASE HELP ME. PRETEND YOU ARE IN A MATH COMPERRITION.

OpenStudy (magbak):

@phi @Math81

OpenStudy (anonymous):

8=6x+6 -6x=6-8 -6x=-2 x=-1/3

OpenStudy (magbak):

show some work please. Like the first one that is good.

OpenStudy (magbak):

Yes thank you SOOOOO MUCH.

OpenStudy (magbak):

I need help with nly one more. PLEASE

OpenStudy (magbak):

Write in simplest form 

OpenStudy (anonymous):

\[\frac{ 2logx+logy }{ 3logz }\]

OpenStudy (magbak):

OK thank you so much.

OpenStudy (phi):

for \[ \log _{2}3+\log _{2}x=3 \] use this rule \[ \log(a) + \log(b)= \log(a\cdot b) \] you get \[ \log_2(3x) = 3\] make these the exponents of 2 \[ 2^{ \log_2(3x) } = 2^3 \] the left side simplifies to 3x \[ 3x = 8 \] solve for x \[ x = \frac{8}{3} \]

OpenStudy (anonymous):

Oh shoot, you're right. I can't believe I missed that

OpenStudy (magbak):

NOw he is correct. I got every single question with you incorrect.

OpenStudy (phi):

this is good until the last line 8=6x+6 -6x=6-8 -6x=-2 x=-1/3 <--- this should be -2/-6 which simplifies to +1/3 (two minuses make a plus)